# Geometrical properties of strained and twisted moiré heterostructures

### v1

@@ I Introduction
A lattice misalignment in stacked two-dimensional (2D) materials leads to moiré superlattices with tunable periodicity [he2021moire, andrei2021marvels]. For example, bilayer graphene with a twist angle $\theta\sim 1.1^{\circ}$ generates a moiré pattern with a period of $\sim 14$ nm, which is termed as magic angle twisted bilayer graphene (TBG) [andrei2020graphene]. Such moiré pattern can cause the appearance of flat bands near zero Fermi energy in the electronic band structure [lopes2007graphene, bistritzer2011moire], which are found to host an amount of topological and interaction-driven many-body states, for example, correlated insulating states [cao2018correlated, xu2020correlated], superconductivity [cao2018unconventional, xia2025superconductivity, guo2025superconductivity], ferroelectricity [zheng2020unconventional], ferromagnetism [sharpe2019emergent], integer and fractional quantum anomalous Hall states [serlin2020intrinsic, xu2023observation, cai2023signatures, zeng2023thermodynamic, park2023observation], phase transitions [li2021continuous] and so on. Similar correlated phenomena have also been observed in other moiré heterostructures, such as twisted trilayer graphene [park2021tunable, kim2022evidence], twisted multilayer graphene [zhang2022promotion], and twisted bilayer $\mathrm{WSe_{2}}$ [wang2020correlated, xia2025superconductivity, guo2025superconductivity].
@@
These rich correlated phenomena typically emerge at moiré scales much larger than the underlying atomic lengths, whereby the moiré acts as a magnifying glass of the underlying geometrical and electronic properties [cosma2014moire]. This makes the moiré superlattices a flexible and tunable system, with plenty of tuning knobs such as twist angle, strain, external field, and substrate. Among these, strain stands out as a powerful tool to modify the geometrical and electronic properties of the stacked lattices [naumis2017electronic, naumis2023mechanical, dai2019strain]. In contrast to twist-only configurations, which can only lead to the same moiré geometry of the underlying lattices, strain-induced deformations can lead to a plethora of different moiré geometries [kogl2023moire, escudero2024designing]. Crucially, due to the moiré magnifying effect, the superlattice pattern can be significantly modified even if the layers are barely deformed [cosma2014moire, escudero2024designing]. As a result, strains can strongly change the moiré size, the superlattice symmetry, and by extension, the electronic properties of the system.
@@
While strain in moiré superlattices is well known to arise naturally during their fabrication [kazmierczak2021strain], recent experimental advancements have opened the possibility of actively induce and control external strains in the system [kapfer2023programming, pena2023moire, ou2025continuous]. A precise control on both twist and strain greatly magnifies the tunability of moiré superlattices, paving the way for potentially novel correlated phenomena. Since the electronic properties of moiré systems depend largely on the electronic modulation due to the moiré potential, it becomes imperative to understand precisely how the geometrical properties of moiré heterostructures depend on the interplay between twist and strain. Among different types of strain, uniaxial heterostrain is the most common type, and has been observed in many experimental samples [huder2018electronic, kerelsky2019maximized, xie2019spectroscopic, mesple2021heterostrain, wang2023unusual]. Recently, under some new well-developed techniques, shear heterostrain is achieved, resulting in deformed moiré patterns [yu2024twist, carrasco2025].
@@
Although recent works have already provide a review of the strain effect and implementation in moiré heterostructures [he2021moire, hou2025strain, zhai2025twistronics, trambly2026electronic], there is yet no detailed review of the geometrical description. Namely, on how does the geometry of the moiré patterns depend on the interplay between twist and strain, and in particular, on the different types of strain configurations that can be engineered in experimental setups. The goal of this review is to cover this gap by offering an overview of the recent advances on the treatment and implementation of strain in moiré heterostructures. To make this work comprehensive, we start by first revisiting the linear elasticity theory, adapted to common in-plane stress and strain in 2D materials. Next, we apply the linear elasticity theory to the twisted and strained moiré systems, and describe how many different moiré geometries can be archived by tuning both the twist and strain. Finally, we discuss recent major breakthroughs in implementing strain in moiré heterostructures, and show experimental examples of particular moiré geometries.
@@ II Linear elasticity in two-dimensional materials
In moiré heterostructures made up of 2D materials (one-atom thickness), the most relevant stress-induced deformations are planar, i.e., in-plane displacements that act on the 2D plane of the material [pereira2009tight]. However, there can also be out-of-plane displacements in the atomic positions (e.g., corrugations). For instance, thermal ripples are well known to appear naturally in graphene at room temperatures [meyer2007structure, fasolino2007intrinsic]. Moreover, out-of-plane corrugations emerge naturally when one stacks two or more 2D materials with a lattice mismatch (induced, e.g., by twist or strain), as the system can reduce its energy by increasing the distance at the energetically costly stacking regimes [uchida2014atomic, nam2017lattice, koshino2018maximally, cazeaux2023relaxation].
In this review, we will simplify the problem by considering only in-plane stresses. This is a safe approximation as our focus is only to describe the geometrical properties of the emergent in-plane moiré patterns. We label the 2D materials plane with the coordinates $x$ and $y$. For self-consistence, and to set up the ground for the following sections, in this section we briefly review the theory of linear elasticity [landau2012theory, gurtin1982introduction, gurtin2010mechanics], adapted to in-plane stress and strain in 2D materials.
@@ II.1 Displacement field, strain and stress
In elasticity theory, the relevant quantity that identifies a deformation of a material is the displacement field $\mathbf{u}$, defined as [landau2012theory, gurtin2010mechanics]
@@
where $\mathbf{r}^{\prime}=\left(x^{\prime},y^{\prime}\right)$ and $\mathbf{r}=\left(x,y\right)$ are the deformed and undeformed position of the material, respectively. In general, the displacement vector could also depend on the time $t$, i.e., $\mathbf{u}\left(\mathbf{r},t\right)$. Here, we simplify the problem and consider time-independent deformations, as is usually the case in two-dimensional materials [vozmediano2010gauge, guinea2012strain, amorim2016novel, si2016strain]. Using that $d\mathbf{r}^{\prime}=d\mathbf{r}+d\mathbf{u}$ and $d\mathbf{u}=\nabla\mathbf{u}d\mathbf{r}$ (where $\nabla$ is the gradient operator), the infinitesimal length $d\ell^{\prime}=\left|d\mathbf{r}^{\prime}\right|$ of the deformed material reads [landau2012theory]
@@
In most practical cases, the deformation of a material is relatively small [amorim2016novel, kazmierczak2021strain, mesple2021heterostrain]. Even in very strong two-dimensional materials like graphene –which can stand up to about 20% strain [lee2008measurement, papageorgiou2017mechanical]– most experimental deformations are at most within the range of $1\%$ and $2\%$ [frisenda2017biaxial, qiao2018twisted, mesple2021heterostrain, kazmierczak2021strain]. We refer to these as small deformations. In such case one can neglect the terms $\mathcal{O}\left(\nabla\mathbf{u}\right)^{2}$ and take [landau2012theory]
@@
The linear theory of elasticity follows the regime of small deformations by keeping all the relevant quantities only up to leading order (linear) in the strain tensor [gurtin2010mechanics]. An important quantity, besides the strain tensor, is the so-called stress tensor $\mathcal{T}$. It is defined so that the tension $\mathbf{t}$ (force per unit length in two-dimensions), at a surface point with normal vector $\mathbf{n}$, is given by the Cauchy’s theorem $\mathbf{t}=\mathcal{T}\mathbf{n}$ [gurtin2010mechanics] (for in-plane stress and strain, the normal vector $\mathbf{n}$ is either of the two in-plane directions, i.e., $\mathbf{n}=\mathbf{e}_{x}$ or $\mathbf{n}=\mathbf{e}_{y}$, or any combination of them). Within the linear theory of elasticity, the stress tensor $\mathcal{T}$ is related to the strain tensor through a generalized Hook’s law [gurtin2010mechanics]
@@
where $\mathbb{C}$ is the fourth-order elasticity tensor (also called stiffness or compliance tensor [pereira2009tight]). Equation (6) gives the stress-strain relation that connects forces to deformations. In practice, the elasticity tensor $\mathbb{C}$ is determined by constitutive equations.
For isotropic materials the relation $\mathcal{T}=\mathbb{C}\mathcal{E}$ becomes a linear isotropic function of the strain tensor; its most general representation reads [gurtin1982introduction]
@@
where $\mu$ and $\lambda$ are known as the Lamé coefficients, and $\mathbb{I}$ is the $2\times 2$ identity matrix. One can then invert Eq. (7) and write the strain tensor in terms of the stress tensor as [gurtin2010mechanics]
@@ II.2 Uniaxial, shear and biaxial strain
In what follows we will limit our discussion to homogeneous (position-independent) strain (for reviews of non-homogeneous strain, see e.g. Refs. [amorim2016novel, naumis2017electronic]). The three experimentally-relevant types of stress in 2D materials are uniaxial, shear, and biaxial, shown in Figure 1.
@@ II.2.1 Uniaxial strain
are the longitudinal strain magnitude and Poisson’s ratio, respectively [landau2012theory, gurtin1982introduction]. A related quantity is the Young’s modulus $E$, given by the ratio between the longitudinal strain and the stress
@@
so the Poisson’s ratio measures the contraction lateral to the direction of the applied stress [gurtin2010mechanics]. The magnitude of such contraction is determined by the material-dependent value of the Poisson’s ratio (e.g., $\nu\sim 0.16$ in graphene [blakslee1970elastic, pereira2009tight]).
If the uniaxial stress acts along a direction $\phi$ with respect to the $x$-axis, then the uniaxial strain becomes [pereira2009tight, bi2019designing]
@@ II.2.2 Shear strain
where $\epsilon_{s}=\sigma_{s}/2\mu$ is the shear strain magnitude. The Lamé coefficient $\mu$ is thus also called the shear modulus [landau2012theory, gurtin2010mechanics]. The shear displacement vector reads $\mathbf{u}=2\epsilon_{s}x\mathbf{e}_{y}$. If the shear acts, in general, perpendicular to a direction $\phi$ (relative to the $x$-axis), the shear strain reads [kogl2023moire, escudero2024designing]
@@ II.3 Deformation of bonds
In the absence of external body forces, a displacement field $\mathbf{u}$ with constant gradient $\nabla\mathbf{u}$ is always a static solution to the equation of motion [landau2012theory, gurtin1982introduction, gurtin2010mechanics]. A constant gradient $\nabla\mathbf{u}$ corresponds to a homogeneous (position-independent) strain tensor [cf. Eq. (5)], as considered in Section II.2. The displacement field $\mathbf{u}$ can then be written in terms of the strain tensor as
@@
In index notation, $u_{i}=\epsilon_{ij}x_{j}$, where repeated indices are to be summed over. Indeed, taking the gradient of $\mathbf{u}=\mathcal{E}\mathbf{r}$, and using the symmetry of the strain tensor, directly leads to Eq. (5). Here we have fixed, without loss of generality, the boundary condition by taking the displacement equal to zero at the origin (i.e., $\mathbf{u}=\mathbf{0}$ when $\mathbf{r}=0$). At small deformations, the deformed bonds (or lengths) of a material then become [pereira2009tight]
@@ III Twisted and strained moiré geometries
Two-dimensional materials stacked in a multilayer configuration, subject to relative twist and strain, create a lattice mismatch between the layers and, consequently, a moiré pattern [moon2013optical, bi2019designing, kogl2023moire, escudero2024designing]. Such structures are thus referred as moiré heterostructures [andrei2021marvels, kennes2021moire]. In this review we will mostly focus on the simplest case of two layers stacked in a bilayer configuration [andrei2020graphene]. To simplify the discussion, we start by considering the simplest (but most common) case of two equal hexagonal lattices. Generalizations to different layers are discussed in subsequent sections.
@@ III.1 Hexagonal homobilayers
We describe the honeycomb lattices by a set of lattice vectors [castro2009electronic]
@@
To describe the geometry of the twisted and strained honeycomb layers, is convenient to consider the effect of twist and strain separately. Specifically, we consider the effect of first twisting and then straining [bi2019designing, sinner2023strain, kogl2023moire, escudero2024designing]. A twist by an angle $\theta$ simply rotates the atomic bonds as $\mathbf{r}\rightarrow R\left(\theta\right)\mathbf{r}$. A further strain on the rotated system will then, according to Eq. (32), transform the atomic positions as $\mathbf{r}^{\prime}=\left(\mathbb{I}+\mathcal{E}\right)R\left(\theta\right)\mathbf{r}$.
@@
where ++ and - refer to the top and bottom layers, respectively. Within the small deformation limit one often keeps only the linear terms in the twist and strain, so that [bi2019designing]
@@
where the symmetric part $\mathcal{E}$ gives the strain effect and the antisymmetric part $\mathcal{R}\left(\theta\right)$ gives the twist effect. In general, the three components of the strain tensor can be decomposed as a mixture of biaxial and shear strains [halbertal2021moire, kogl2023moire]
@@
where $\cos\phi=\epsilon_{xy}/\epsilon_{s}$. In this parameterization the uniaxial strain becomes a mixture of biaxial and shear strain [kogl2023moire] (more details in Section III.1.4).
A symmetric bilayer configuration is to twist and strain the two layers with the same magnitude but in opposite direction [bi2019designing, sinner2023strain, escudero2024designing], so that $\theta$$_{\pm}=\pm\theta/2$ and $\mathcal{E}_{\pm}=\pm\mathcal{E}/2$. But there are, of course, many other possibilities. For instance, one layer could be only rotated while the other layer is only strained [kogl2023moire]. For conciseness and simplicity, in this section we will consider the symmetric configuration. The presented analysis is, however, general and does not depend on the particular twist and strain in each layer.
In general, for arbitrary twist and strain, the moiré pattern is not commensurate [bi2019designing, escudero2024designing]. Commensuration only occurs at particular twist and strain for which one can find a superlattice vector common to both layers [shallcross2010electronic, moon2013optical, shi2020large, wang2025]. Here we shall not pursue the problem of finding commensurate solutions under both twist and strain. As we will see, the geometrical properties of the moiré pattern can be well defined and studied even if the system is technically incommensurate [koshino2015interlayer]. In fact, in the relevant regime of moiré scales much larger than the atomic lengths, the electronic properties become insensitive to the commensurability of the system [bistritzer2011moire]. Effective continuum models, widely employed to study the electronic properties of large moiré heterostructures, can be well defined even if the system is incommensurate [koshino2015interlayer, bi2019designing]. The commensuration is, however, fundamental when one needs to compute the band structure by more realistic methods, such as atomistic tight-binding [suarez2010flat, long2022atomistic, wang2025] or density functional theory (DFT) simulations [cantele2020structural, leconte2022relaxation].
The geometrical properties of the moiré pattern are more easily described by working in reciprocal space [bi2019designing, escudero2024designing]. The twisted and strained reciprocal lattice vectors $\mathbf{b}_{i,\pm}$ in each layer follow from Eq. (36) and the condition $\mathbf{a}_{i,\pm}\cdot\mathbf{b}_{j,\pm}=2\pi\delta_{ij}$, which implies
@@
The last approximation holds within the small deformation regime [bi2019designing]. The moiré vectors of the moiré pattern can be formally defined as the set of vectors $\mathbf{G}_{M}$ that satisfy
@@
where $\mathbf{R}_{M}$ are the minimal union of the lattice vectors $\mathbf{R}_{\pm}=n_{1}\mathbf{a}_{1,\pm}+n_{2}\mathbf{a}_{2,\pm}$ in each layer (with $n_{1}$ and $n_{2}$ integers). In a commensurate structure $\mathbf{R}_{M}$ reduce to the superlattice vectors at which $\mathbf{R}_{M}=\mathbf{R}_{\pm}$ [shallcross2010electronic], but in an incommensurate structure $\mathbf{R}_{M}$ is generally given by the lattice points at which the difference between $\mathbf{R}_{+}$ and $\mathbf{R}_{-}$ is minimum, which always occurs around the AA stacking regimes [moon2013optical]. The primitive moiré vectors $\mathbf{G}_{i}$ ($i=1,2$) are the two smallest vectors $\mathbf{G}_{M}$ from which all others can be obtained by translation. In the case of only a twist, the primitive moiré vectors are given by the difference between the reciprocal lattice vectors [moon2013optical, koshino2015interlayer, koshino2018maximally]
@@
However, this is not necessarily the case in the presence of strain. Indeed, the deformation of each lattice may be such that other combination between $\mathbf{b}_{i,+}$ and $\mathbf{b}_{i,-}$ may yield a smaller moiré vector [escudero2024designing] (see Figure 2). Accounting for these different constructions of the moiré vectors is essential to elucidate the wide range of moiré patterns that can be formed by strain. The analysis can be simplified by noting that the different geometrical constructions of the primitive moiré vectors must be related by the hexagonal symmetry of the underlying lattices. This means that it is sufficient to study, as usually done, only the geometrical properties of the moiré vectors given by Eq. (45), keeping in mind that all other constructions are related by $60^{\circ}$ translations of the strain parameters [escudero2024designing].
Figure 2: Strain-dependent construction of moiré vectors. The left panels shows the deformed reciprocal vectors $\mathbf{b}_{i,\pm}$ in two honeycomb lattices subject to a twist $\theta=5^{\circ}$ and uniaxial strain with magnitude $\epsilon=5\%$ and direction $\phi=60^{\circ}$. Taking the difference between the deformed reciprocal vectors does not yield, in this case, the smallest moiré vectors $\mathbf{G}_{i}$. A smaller moiré vector $\mathbf{G}^{\prime}_{2}=\mathbf{G}_{1}+\mathbf{G}_{2}$ is obtained by taking the difference between the deformed reciprocal vectors $\mathbf{b}^{\prime}_{2,\pm}=\mathbf{b}_{1,\pm}+\mathbf{b}_{2,\pm}$. The right panel shows the superlattice spanned by the moiré vectors. Adapted under the terms of the $C$ BY license from Ref. [escudero2024designing]. Copyright (2024) by the American Physical Society.
For stacking configurations made of two homobilayers, the moiré vectors given by Eq. (45) are determined by a unique transformation $\mathbf{T}$ acting on the reciprocal vectors [sinner2023strain], $\mathbf{G}_{i}=\mathbf{T}\mathbf{b}_{i}$, where (for the symmetrical configuration)

### v2

@@ I Introduction
A lattice misalignment in stacked two-dimensional (2D) materials leads to moiré superlattices with tunable periodicity [47, 3]. For example, bilayer graphene with a twist angle $\theta\sim 1.1^{\circ}$ generates a moiré pattern with a period of $\sim 14$ nm, which is termed as magic angle twisted bilayer graphene (TBG) [4]. Such moiré pattern can cause the appearance of flat bands near zero Fermi energy in the electronic band structure [78, 9], which are found to host an amount of topological and interaction-driven many-body states, for example, correlated insulating states [15, 123], superconductivity [16, 120, 42], ferroelectricity [136], ferromagnetism [103], integer and fractional quantum anomalous Hall states [100, 122, 13, 129, 91], phase transitions [72] and so on. Similar correlated phenomena have also been observed in other moiré heterostructures, such as twisted trilayer graphene [92, 59], twisted multilayer graphene [133], and twisted bilayer $\mathrm{WSe_{2}}$ [117, 120, 42].
@@
These rich correlated phenomena typically emerge at moiré scales much larger than the underlying atomic lengths, whereby the moiré acts as a magnifying glass of the underlying geometrical and electronic properties [27]. This makes the moiré superlattices a flexible and tunable system, with plenty of tuning knobs such as twist angle, strain, external field, and substrate. Among these, strain stands out as a powerful tool to modify the geometrical and electronic properties of the stacked lattices [86, 87, 28]. In contrast to twist-only configurations, which can only lead to the same moiré geometry of the underlying lattices, strain-induced deformations can lead to a plethora of different moiré geometries [60, 34]. Crucially, due to the moiré magnifying effect, the superlattice pattern can be significantly modified even if the layers are barely deformed [27, 34]. As a result, strains can strongly change the moiré size, the superlattice symmetry, and by extension, the electronic properties of the system.
@@
While strain in moiré superlattices is well known to arise naturally during their fabrication [56], recent experimental advancements have opened the possibility of actively induce and control external strains in the system [55, 94, 88]. A precise control on both twist and strain greatly magnifies the tunability of moiré superlattices, paving the way for potentially novel correlated phenomena. Since the electronic properties of moiré systems depend largely on the electronic modulation due to the moiré potential, it becomes imperative to understand precisely how the geometrical properties of moiré heterostructures depend on the interplay between twist and strain. Among different types of strain, uniaxial heterostrain is the most common type, and has been observed in many experimental samples [51, 58, 121, 81, 119]. Recently, under some new well-developed techniques, shear heterostrain is achieved, resulting in deformed moiré patterns [127, 19].
@@
Although recent works have already provide a review of the strain effect and implementation in moiré heterostructures [47, 50, 130, 112], there is yet no detailed review of the geometrical description. Namely, on how does the geometry of the moiré patterns depend on the interplay between twist and strain, and in particular, on the different types of strain configurations that can be engineered in experimental setups. The goal of this review is to cover this gap by offering an overview of the recent advances on the treatment and implementation of strain in moiré heterostructures. To make this work comprehensive, we start by first revisiting the linear elasticity theory, adapted to common in-plane stress and strain in 2D materials. Next, we apply the linear elasticity theory to the twisted and strained moiré systems, and describe how many different moiré geometries can be archived by tuning both the twist and strain. Finally, we discuss recent major breakthroughs in implementing strain in moiré heterostructures, and show experimental examples of particular moiré geometries.
@@ II Linear elasticity in two-dimensional materials
In moiré heterostructures made up of 2D materials (one-atom thickness), the most relevant stress-induced deformations are planar, i.e., in-plane displacements that act on the 2D plane of the material [95]. However, there can also be out-of-plane displacements in the atomic positions (e.g., corrugations). For instance, thermal ripples are well known to appear naturally in graphene at room temperatures [83, 37]. Moreover, out-of-plane corrugations emerge naturally when one stacks two or more 2D materials with a lattice mismatch (induced, e.g., by twist or strain), as the system can reduce its energy by increasing the distance at the energetically costly stacking regimes [113, 85, 61, 22].
In this review, we will simplify the problem by considering only in-plane stresses. This is a safe approximation as our focus is only to describe the geometrical properties of the emergent in-plane moiré patterns. We label the 2D materials plane with the coordinates $x$ and $y$. For self-consistence, and to set up the ground for the following sections, in this section we briefly review the theory of linear elasticity [65, 44, 43], adapted to in-plane stress and strain in 2D materials.
@@ II.1 Displacement field, strain and stress
In elasticity theory, the relevant quantity that identifies a deformation of a material is the displacement field $\mathbf{u}$, defined as [65, 43]
@@
where $\mathbf{r}^{\prime}=\left(x^{\prime},y^{\prime}\right)$ and $\mathbf{r}=\left(x,y\right)$ are the deformed and undeformed position of the material, respectively. In general, the displacement vector could also depend on the time $t$, i.e., $\mathbf{u}\left(\mathbf{r},t\right)$. Here, we simplify the problem and consider time-independent deformations, as is usually the case in two-dimensional materials [116, 40, 2, 106]. Using that $d\mathbf{r}^{\prime}=d\mathbf{r}+d\mathbf{u}$ and $d\mathbf{u}=\nabla\mathbf{u}d\mathbf{r}$ (where $\nabla$ is the gradient operator), the infinitesimal length $d\ell^{\prime}=\left|d\mathbf{r}^{\prime}\right|$ of the deformed material reads [65]
@@
In most practical cases, the deformation of a material is relatively small [2, 56, 81]. Even in very strong two-dimensional materials like graphene –which can stand up to about 20% strain [67, 90]– most experimental deformations are at most within the range of $1\%$ and $2\%$ [38, 97, 81, 56]. We refer to these as small deformations. In such case one can neglect the terms $\mathcal{O}\left(\nabla\mathbf{u}\right)^{2}$ and take [65]
@@
The linear theory of elasticity follows the regime of small deformations by keeping all the relevant quantities only up to leading order (linear) in the strain tensor [43]. An important quantity, besides the strain tensor, is the so-called stress tensor $\mathcal{T}$. It is defined so that the tension $\mathbf{t}$ (force per unit length in two-dimensions), at a surface point with normal vector $\mathbf{n}$, is given by the Cauchy’s theorem $\mathbf{t}=\mathcal{T}\mathbf{n}$ [43] (for in-plane stress and strain, the normal vector $\mathbf{n}$ is either of the two in-plane directions, i.e., $\mathbf{n}=\mathbf{e}_{x}$ or $\mathbf{n}=\mathbf{e}_{y}$, or any combination of them). Within the linear theory of elasticity, the stress tensor $\mathcal{T}$ is related to the strain tensor through a generalized Hook’s law [43]
@@
where $\mathbb{C}$ is the fourth-order elasticity tensor (also called stiffness or compliance tensor [95]). Equation (6) gives the stress-strain relation that connects forces to deformations. In practice, the elasticity tensor $\mathbb{C}$ is determined by constitutive equations.
For isotropic materials the relation $\mathcal{T}=\mathbb{C}\mathcal{E}$ becomes a linear isotropic function of the strain tensor; its most general representation reads [44]
@@
where $\mu$ and $\lambda$ are known as the Lamé coefficients, and $\mathbb{I}$ is the $2\times 2$ identity matrix. One can then invert Eq. (7) and write the strain tensor in terms of the stress tensor as [43]
@@ II.2 Uniaxial, shear and biaxial strain
In what follows we will limit our discussion to homogeneous (position-independent) strain (for reviews of non-homogeneous strain, see e.g. Refs. [2, 86]). The three experimentally-relevant types of stress in 2D materials are uniaxial, shear, and biaxial, shown in Figure 1.
@@ II.2.1 Uniaxial strain
are the longitudinal strain magnitude and Poisson’s ratio, respectively [65, 44]. A related quantity is the Young’s modulus $E$, given by the ratio between the longitudinal strain and the stress
@@
so the Poisson’s ratio measures the contraction lateral to the direction of the applied stress [43]. The magnitude of such contraction is determined by the material-dependent value of the Poisson’s ratio (e.g., $\nu\sim 0.16$ in graphene [10, 95]).
If the uniaxial stress acts along a direction $\phi$ with respect to the $x$-axis, then the uniaxial strain becomes [95, 8]
@@ II.2.2 Shear strain
where $\epsilon_{s}=\sigma_{s}/2\mu$ is the shear strain magnitude. The Lamé coefficient $\mu$ is thus also called the shear modulus [65, 43]. The shear displacement vector reads $\mathbf{u}=2\epsilon_{s}x\mathbf{e}_{y}$. If the shear acts, in general, perpendicular to a direction $\phi$ (relative to the $x$-axis), the shear strain reads [60, 34]
@@ II.3 Deformation of bonds
In the absence of external body forces, a displacement field $\mathbf{u}$ with constant gradient $\nabla\mathbf{u}$ is always a static solution to the equation of motion [65, 44, 43]. A constant gradient $\nabla\mathbf{u}$ corresponds to a homogeneous (position-independent) strain tensor [cf. Eq. (5)], as considered in Section II.2. The displacement field $\mathbf{u}$ can then be written in terms of the strain tensor as
@@
In index notation, $u_{i}=\epsilon_{ij}x_{j}$, where repeated indices are to be summed over. Indeed, taking the gradient of $\mathbf{u}=\mathcal{E}\mathbf{r}$, and using the symmetry of the strain tensor, directly leads to Eq. (5). Here we have fixed, without loss of generality, the boundary condition by taking the displacement equal to zero at the origin (i.e., $\mathbf{u}=\mathbf{0}$ when $\mathbf{r}=0$). At small deformations, the deformed bonds (or lengths) of a material then become [95]
@@ III Twisted and strained moiré geometries
Two-dimensional materials stacked in a multilayer configuration, subject to relative twist and strain, create a lattice mismatch between the layers and, consequently, a moiré pattern [84, 8, 60, 34]. Such structures are thus referred as moiré heterostructures [3, 57]. In this review we will mostly focus on the simplest case of two layers stacked in a bilayer configuration [4]. To simplify the discussion, we start by considering the simplest (but most common) case of two equal hexagonal lattices. Generalizations to different layers are discussed in subsequent sections.
@@ III.1 Hexagonal homobilayers
We describe the honeycomb lattices by a set of lattice vectors [21]
@@
To describe the geometry of the twisted and strained honeycomb layers, is convenient to consider the effect of twist and strain separately. Specifically, we consider the effect of first twisting and then straining [8, 107, 60, 34]. A twist by an angle $\theta$ simply rotates the atomic bonds as $\mathbf{r}\rightarrow R\left(\theta\right)\mathbf{r}$. A further strain on the rotated system will then, according to Eq. (32), transform the atomic positions as $\mathbf{r}^{\prime}=\left(\mathbb{I}+\mathcal{E}\right)R\left(\theta\right)\mathbf{r}$.
@@
where ++ and - refer to the top and bottom layers, respectively. Within the small deformation limit one often keeps only the linear terms in the twist and strain, so that [8]
@@
where the symmetric part $\mathcal{E}$ gives the strain effect and the antisymmetric part $\mathcal{R}\left(\theta\right)$ gives the twist effect. In general, the three components of the strain tensor can be decomposed as a mixture of biaxial and shear strains [45, 60]
@@
where $\cos\phi=\epsilon_{xy}/\epsilon_{s}$. In this parameterization the uniaxial strain becomes a mixture of biaxial and shear strain [60] (more details in Section III.1.4).
A symmetric bilayer configuration is to twist and strain the two layers with the same magnitude but in opposite direction [8, 107, 34], so that $\theta$$_{\pm}=\pm\theta/2$ and $\mathcal{E}_{\pm}=\pm\mathcal{E}/2$. But there are, of course, many other possibilities. For instance, one layer could be only rotated while the other layer is only strained [60]. For conciseness and simplicity, in this section we will consider the symmetric configuration. The presented analysis is, however, general and does not depend on the particular twist and strain in each layer.
In general, for arbitrary twist and strain, the moiré pattern is not commensurate [8, 34]. Commensuration only occurs at particular twist and strain for which one can find a superlattice vector common to both layers [102, 84, 104, 35]. Here we shall not pursue the problem of finding commensurate solutions, and consider arbitrary twist and strain regardless of whether they correspond to a commensurate or incommensurate moiré. As we will see, the geometrical properties of the moiré pattern can be well defined and studied even if the system is technically incommensurate [62]. In fact, in the relevant regime of moiré scales much larger than the atomic lengths, the electronic properties become insensitive to the commensurability of the system [9]. Effective continuum models, widely employed to study the electronic properties of large moiré heterostructures, can be well defined even if the system is incommensurate [62, 8]. The commensuration is, however, fundamental when one needs to compute the band structure by more realistic methods, such as atomistic tight-binding [109, 76, 35] or density functional theory (DFT) simulations [14, 66].
The geometrical properties of the moiré pattern are more easily described by working in reciprocal space [8, 34]. The twisted and strained reciprocal lattice vectors $\mathbf{b}_{i,\pm}$ in each layer follow from Eq. (36) and the condition $\mathbf{a}_{i,\pm}\cdot\mathbf{b}_{j,\pm}=2\pi\delta_{ij}$, which implies
@@
The last approximation holds within the small deformation regime [8]. The moiré vectors of the moiré pattern can be formally defined as the set of vectors $\mathbf{G}_{M}$ that satisfy
@@
where $\mathbf{R}_{M}$ are the minimal union of the lattice vectors $\mathbf{R}_{\pm}=n_{1}\mathbf{a}_{1,\pm}+n_{2}\mathbf{a}_{2,\pm}$ in each layer (with $n_{1}$ and $n_{2}$ integers). In a commensurate structure $\mathbf{R}_{M}$ reduce to the superlattice vectors at which $\mathbf{R}_{M}=\mathbf{R}_{\pm}$ [102], but in an incommensurate structure $\mathbf{R}_{M}$ is generally given by the lattice points at which the difference between $\mathbf{R}_{+}$ and $\mathbf{R}_{-}$ is minimum, which always occurs around the AA stacking regimes [84]. The primitive moiré vectors $\mathbf{G}_{i}$ ($i=1,2$) are the two smallest vectors $\mathbf{G}_{M}$ from which all others can be obtained by translation. In the case of only a twist, the primitive moiré vectors are given by the difference between the reciprocal lattice vectors [84, 62, 61]
@@
However, this is not necessarily the case in the presence of strain. Indeed, the deformation of each lattice may be such that other combination between $\mathbf{b}_{i,+}$ and $\mathbf{b}_{i,-}$ may yield a smaller moiré vector [34] (see Figure 2). Accounting for these different constructions of the moiré vectors is essential to elucidate the wide range of moiré patterns that can be formed by strain. The analysis can be simplified by noting that the different geometrical constructions of the primitive moiré vectors must be related by the hexagonal symmetry of the underlying lattices. This means that it is sufficient to study, as usually done, only the geometrical properties of the moiré vectors given by Eq. (45), keeping in mind that all other constructions are related by $60^{\circ}$ translations of the strain parameters [34].
Figure 2: Strain-dependent construction of moiré vectors. The left panels shows the deformed reciprocal vectors $\mathbf{b}_{i,\pm}$ in two honeycomb lattices subject to a twist $\theta=5^{\circ}$ and uniaxial strain with magnitude $\epsilon=5\%$ and direction $\phi=60^{\circ}$. Taking the difference between the deformed reciprocal vectors does not yield, in this case, the smallest moiré vectors $\mathbf{G}_{i}$. A smaller moiré vector $\mathbf{G}^{\prime}_{2}=\mathbf{G}_{1}+\mathbf{G}_{2}$ is obtained by taking the difference between the deformed reciprocal vectors $\mathbf{b}^{\prime}_{2,\pm}=\mathbf{b}_{1,\pm}+\mathbf{b}_{2,\pm}$. The right panel shows the superlattice spanned by the moiré vectors. Adapted under the terms of the $C$ BY license from Ref. [34]. Copyright (2024) by the American Physical Society.
For stacking configurations made of two homobilayers, the moiré vectors given by Eq. (45) are determined by a unique transformation $\mathbf{T}$ acting on the reciprocal vectors [107], $\mathbf{G}_{i}=\mathbf{T}\mathbf{b}_{i}$, where (for the symmetrical configuration)